OpenStudy (anonymous):

he product of two consecutive negative integers is 1122. What are the numbers he product of two consecutive negative integers is 1122. What are the numbers @Mathematics

OpenStudy (diyadiya):

x(x+1) = 1122 x^2 + x = 1122 x^2 + x - 1122 = 0

OpenStudy (diyadiya):

(x-33)(x+34)

OpenStudy (diyadiya):

two numbers can be taken as -34 and -33

OpenStudy (anonymous):

\[ n(n+1)= 1122 \Rightarrow = n= 33,-34 \] so the numbers are -34 and -33.